# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v2

@@ III Quantization of the regularized quasinormal modes
We now introduce a quantization scheme for the regularized QNMs of arbitrary, 3D, linear, dissipative (with possible gain components) and dispersive nanoresonators, as defined in § II. In our approach, we will quantize the eigenmodes of the PML-truncated domain in Eqs. (II.1) and (II.1). As already mentioned, this set of eigenmodes includes QNM-like and PML modes, which following our earlier remarks, shall simply be referred to as modes without further differentiation.
@@
Similar to previous work Franke et al. (2019, 2020a); Franke, Ren, and Hughes (2022), our approach is predicated on the formalism of mQED, which offers a phenomenological Green’s function quantization scheme for electromagnetic fields in the presence of spatially inhomogeneous, lossy and dispersive media Gruner and Welsch (1996); Dung, Knöll, and Welsch (1998); Scheel, Knöll, and Welsch (1998); Dung, Knöll, and Welsch (2000); Raabe, Scheel, and Welsch (2007), later generalized for systems with gain Matloob et al. (1997); Raabe and Welsch (2008). However, in order to facilitate a rigorous application of the theory, a brief discussion concerning some important prerequisites is warranted. The complex permittivity $\varepsilon(\bm{r},\omega)=\varepsilon_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\varepsilon_{\mathrm{i}}(\bm{r},\omega)$ and inverse permeability $\mu^{-1}(\bm{r},\omega)=\kappa(\bm{r},\omega)=\kappa_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\kappa_{\mathrm{i}}(\bm{r},\omega)$ are required to satisfy the Kramers-Kronig relations (i.e., they are analytic in the closed, upper-half, complex $\omega$ plane and $\zeta(\bm{r},\omega)\rightarrow 1$ for $|\omega|\rightarrow\infty$ ($\zeta\in\{\varepsilon,\kappa\}$). This is not really a requirement of the quantum theory, but arises from the physical necessity of causality. Furthermore, they should be reciprocal, $\zeta(\bm{r},\omega)=\zeta^{\mathrm{T}}(\bm{r},\omega)$, and satisfy the reflection symmetry condition $\zeta_{ij}(\bm{r},-\omega^{*})=\zeta^{*}_{ij}(\bm{r},\omega)$, which guarantees their reality in the time domain. These properties ensure that the fundamental QED commutation relations between the electric and magnetic field operators are maintained. If the material parameters $\zeta(\bm{r},\omega)$ of the system are reciprocal, it readily follows from Eq. (10) that the PML-transformed parameters $\tilde{\zeta}(\bm{r},\omega)$ are as well. It should be noted that there have been generalizations of mQED for the cases of non-reciprocal materials and materials with magnetoelectric crossing effects Buhmann, Butcher, and Scheel (2012), which are, however, beyond the scope of this work.
@@
When incorporating dispersive or non-dispersive PMLs in the quantization, two problems arise. For dispersive PMLs, it is easy to show from Eq. (10) and the Jacobian in Eq. (9) that the PML-transformed parameters $\tilde{\zeta}(\bm{r},\omega)$ satisfy reflection symmetry, which preserves the important property that modes with $\Re(\tilde{\omega}_{n})\neq 0$ always appear in pairs $(\tilde{\bm{E}}_{\mu},\tilde{\omega}_{\mu})$, $(\tilde{\bm{E}}^{*}_{n},-\tilde{\omega}^{*}_{n})$ and that all the modes have a non-positive imaginary part $\Im(\tilde{\omega}_{n})\leq 0$. Due to the $\omega$-dependence of the coordinate transformation, the material parameters have a pole at $\omega=0$ and consequently do not satisfy the Kramers-Kronig relations, (i.e., they do not correspond to a causal medium). To resolve this issue, we could construct causal PMLs Kuzuoglu and Mittra (2002) by making the replacement $\omega\rightarrow\omega+\mathrm{i}\eta$ in Eq. (6) with a small $\eta>0$, thereby shifting the pole into the lower, complex $\omega$ half-plane. Note that these causal PMLs also work, in principle, in the static case $\omega=0$, due to the finite values of the transformed material parameters at this point Kuzuoglu and Mittra (2002), whereas the usual PMLs fail and cannot be used to reproduce static modes Besbes and Sauvan (2022). However, we are only interested in the fields in the inner, unmapped domain $\Omega\setminus\Omega_{\mathrm{PML}}$, and by using the properties of the Green’s function we will show that the correct commutator is maintained within this region even if it is not in the PML region $\Omega_{\mathrm{PML}}$, so we expect that the use of PMLs as defined in Eq. (6) should suffice in most cases where static modes can be neglected. When using non-dispersive PMLs, the PML-transformed parameters no longer satisfy reflection symmetry, and since they do not afford the damping of waves with non-positive frequencies, the QNMs with $\Re(\tilde{\omega}_{n})\leq 0$ are not correctly recovered and PML modes with $\Im(\tilde{\omega}_{n})>0$ occur, which would grow exponentially in time Sauvan et al. (2022). In practice, one is often interested in the spectral response of a system only within a finite frequency interval $\Delta\omega$, in which case non-dispersive PMLs could be used provided that the contributions of the modes with $\Re(\tilde{\omega}_{n})\leq 0$ are negligible in $\Delta\omega$. With the same argument, one can also justify the use of material parameters $\zeta(\bm{r},\omega)$ or PMLs that do not satisfy the Kramers-Kronig relations (e.g., non-dispersive and complex), provided that the true, physical parameters are sufficiently well approximated within $\Delta\omega$.
@@
×$\nabla\times\tilde{\kappa}\nabla\times G(\bm{r},\bm{r}^{\prime},\omega)-\frac{\omega^{2}}{c^{2}}\tilde{\varepsilon}G(\bm{r},\bm{r}^{\prime},\omega)=\bm{I}\delta(\bm{r}-\bm{r}^{\prime})\,.$ (21)
If the system exhibits loss only, then the Green’s function is analytic in the upper $\omega$ half-plane and has a set of discrete poles (corresponding to the QNMs), as well as possible branch cuts in the lower half-plane in the case of non-trivial background geometries. The introduction of gain components in the system shifts up the poles in the complex $\omega$ plane and a prerequisite to apply the theory for such active systems is that all the poles remain located in the lower half-plane. Again, this is not so much a fundamental limitation of the quantum theory but is, rather, a constraint on its range of validity for systems with linear amplification Raabe and Welsch (2008). The Green’s function thus also satisfies the Kramers-Kronig relations, and besides that, reflection symmetry and the reciprocity $G(\bm{r},\bm{r}^{\prime},\omega)=G^{\mathrm{T}}(\bm{r}^{\prime},\bm{r},\omega)$.
@@
The second ingredient that we need to formulate the modal quantum theory is an expansion of the Green’s function in terms of the modes in Eq. (II.1). By using a residue-decomposition approach, one can derive the expansion
with $A_{n}=\varepsilon_{0}c^{2}/\omega(\tilde{\omega}_{n}-\omega)$ or $A_{n}=\varepsilon_{0}c^{2}/\tilde{\omega}_{n}(\tilde{\omega}_{n}-\omega)$, where the non-uniqueness of the expansion stems from the overcompleteness of the set of modes Lalanne et al. (2018).
@@
×$\nabla\times\tilde{\kappa}\nabla\times\hat{\bm{E}}(\bm{r},\omega)-\frac{\omega^{2}}{c^{2}}\tilde{\varepsilon}\hat{\bm{E}}(\bm{r},\omega)=i\omega\mu_{0}\hat{\bm{j}}_{\mathrm{F}}(\bm{r},\omega)\,.$ (24)
@@
+×$+\nabla\times\bigg{(}\sqrt{-\frac{\hbar}{\pi\mu_{0}}\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega)}\hat{f}_{\mathrm{m},j}(\bm{r},\omega)\bm{e}_{j}\bigg{)}\,,$
in which we have used the Einstein summation convention and $\hat{f}_{\lambda,j}(\bm{r},\omega)$ ($\hat{f}^{\dagger}_{\lambda,j}(\bm{r},\omega)$) are bosonic annihilation (creation) operators of the joint (polaritonic) excitations of the field-dressed medium. The latter satisfy the commutation relationDung, Knöll, and Welsch (1998)
@@
The noise current defined in Eq. (25) stands in accordance with the fluctuation-dissipation theorem and ensures preservation of the fundamental commutation relations of the electric and magnetic fields. These formulae can also be derived from the canonical quantization of a microscopic model in which the electric and magnetic material responses are described using separate oscillator fields Philbin (2010). From this perspective, it is clear that the Fock space corresponding to these bosonic variables describes not only the electromagnetic degrees of freedom, but also those of the material excitations.
@@
+×$+\nabla\times\bigg{(}\sqrt{\frac{\hbar}{\pi\mu_{0}}|\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega)|}\bm{e}_{j}\bigg{[}\Theta(-\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega))\hat{f}_{\mathrm{m},j}(\bm{r},\omega)$
@@
[…]
@@
Note that excitations in the gain regions have a negative-energy contribution to the Hamiltonian. These expressions can again be derived from a microscopic model using inverted oscillator fields in the gain regions Amooghorban et al. (2011). However, it is worth underlining that the expressions in Eqs. (25) and (III) cannot be derived from a generalized phenomenological quantization scheme that allows for general non-local material responses described by a conductivity tensor, whenever the permittivity and permeability are spatially inhomogeneous Raabe and Welsch (2008); Raabe, Scheel, and Welsch (2007). Since these expressions can be motivated from a microscopic model and provide analytic results for arbitrary magnetodielectric media, we will make use of them in this paper.
@@
$=\frac{\hbar\mu_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\int\mathrm{d}^{3}s$
@@
(×$-(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg{]}$
@@
where the latter equation shows that the commutator is not affected by the PMLs for $\bm{r},\bm{r}^{\prime}\in\Omega\setminus\Omega_{\mathrm{PML}}$. By invoking the first equality and the Green’s function expansion in Eq. (23), we can expand the commutator in terms of the modes:
@@
$P_{nn^{\prime}}=\frac{\hbar\mu_{0}}{\pi}$ $\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\int\mathrm{d}^{3}s\,A_{n}(\omega)A_{n^{\prime}}^{*}(\omega)$ (33)
[…]
(×$-(\nabla\times\tilde{\bm{E}}_{n}(\bm{s}))\tilde{\kappa}_{\mathrm{i}}(\bm{s},\omega)(\nabla\times\tilde{\bm{E}}^{*}_{n^{\prime}}(\bm{s}))\bigg{]}\,.$
@@
in which the modal operators can be expressed through the continuous bosonic variables:
$\hat{\xi}^{(\dagger)}_{n,\mathrm{e},\mathrm{L/G}}=\mathrm{i}\mu_{0}\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}A_{n}(\omega)\int\mathrm{d}^{3}r\,\theta(\pm\tilde{\varepsilon}_{\mathrm{i},j}(\bm{r},\omega))$
@@
$\hat{\xi}^{(\dagger)}_{n,\mathrm{m},\mathrm{L/G}}=\mathrm{i}\mu_{0}\int_{0}^{\infty}\mathrm{d}\omega\,\omega A_{n}(\omega)\int\mathrm{d}^{3}r\,\theta(\mp\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega))$
@@
Note that the hermitian conjugates of the gain operators $\hat{\xi}_{n,\mathrm{e/m},\mathrm{G}}$ have to be used in the expansion of the positive-frequency part of the electric field, in order to reproduce the correct commutation relation as given in Eq. (32). Although they are hermitian conjugates of modal operators, they still behave like superpositions of positive-frequency excitations ($\propto\mathrm{e}^{-\mathrm{i}\omega t}$), as can be seen from Eqs. (III) and (36), and the Heisenberg equations for the continuous variables $\hat{f}_{\mathrm{e/m},j}(\bm{r},\omega)$ in the gain regions with the Hamiltonian of Eq. (29). This is analogous to the approach in Ref. Franke, Ren, and Hughes (2022). In our case, however, we have included the PMLs into the quantization, which always exhibit gain. In the presence of gain materials, the vacuum correlation function of the electric field would be given by
@@
$=\frac{\hbar\mu_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\int\mathrm{d}^{3}s$
@@
(×$-(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}|\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)|(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg{]}\ .$
Due to the moduli of the material parameters in this formula, the identity in Eq. (III) cannot be used to express the correlation function in terms of the imaginary part of the Green’s function, which must be taken into account when, for example, calculating QE decay rates Franke et al. (2021) or Casimir(-Polder) forces Sambale et al. (2010). In our approach, this implies that the presence of gain components in the PMLs would affect the correlation function due to induced fluctuations and thus the quantum dynamics, as they would, for example, affect the spontaneous decay of a QE. This means that in order to obtain the correct correlation function of the real system without contributions from the PMLs, we have to eliminate these fluctuations attributable to the gain components of the PMLs from our modal theory. To achieve this, we redefine the gain and loss operators in Eqs. (III) and (36) by including the gain components of the PML region $\Omega_{\mathrm{PML}}$ in the loss operators:
@@
(×$(\nabla\times\bm{E}_{n}(\bm{r}))_{j}(\nabla\times\bm{E}_{n^{\prime}}^{*}(\bm{r}))_{j}$
@@
(×$(\nabla\times\bm{E}_{n}(\bm{r}))_{j}(\nabla\times\bm{E}_{n^{\prime}}^{*}(\bm{r}))_{j}\bigg{]}$
@@
(×$(\nabla\times\bm{E}^{*}_{n}(\bm{r}))_{j}(\nabla\times\bm{E}_{n^{\prime}}(\bm{r}))_{j}$ (45)
All the other commutators vanish due to the bosonic commutation relations of the variables $\hat{\bm{f}}_{\lambda}(\bm{r},\omega)$ and the distinct loss/gain regions. Since the commutation relations of these modal operators are non-bosonic, we have to apply symmetrization transformations separately to the electric/magnetic and loss/gain components in order to ascertain bosonic modal operators Franke et al. (2019); Löwdin (1950):
@@
From the definitions of the modal operators in Eqs. (III)-(III), it is apparent that the loss operators do not have the same vacuum state |$\ket{0}$ as the continuum operators $\hat{\bm{f}}_{\lambda}(\bm{r},\omega)$, as we have included the gain part of the PMLs in the loss operators ($\hat{a}_{n,\mathrm{e/m},\mathrm{L}}\ket{0}\neq 0$). Nevertheless, we adopt it as the vacuum state of the modal operators ($\hat{a}_{n,\mathrm{e/m},\mathrm{L}}\ket{0}=0$) and use it to construct a modal bosonic Fock space $\mathcal{F}$, in which the $k$-photon Fock states can be constructed from |$\ket{0}$ in the usual manner:
|$\ket{k_{n,\lambda,X}}=\frac{1}{\sqrt{k_{n,\lambda,X}!}}\left(\hat{a}^{\dagger}_{n,\lambda,X}\right)^{k}\ket{0}\ .$ (51)
@@
+(×$+(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}|\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)|(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}$

### v3

@@ III Quantization of the Regularized Quasinormal Modes
We now introduce a quantization scheme for the regularized QNMs of arbitrary, 3D, linear, dissipative (with possible gain components) and dispersive nanoresonators, as defined in § II. In our approach, we will quantize the eigenmodes of the PML-truncated domain in Eqs. (II) and (II). As already mentioned, this set of eigenmodes includes QNM-like and PML modes, which following our earlier remarks, shall simply be referred to as modes without further differentiation.
@@
Similar to previous work franke2019quantization; franke2020nonlinear; franke2022coupled, our approach is predicated on the formalism of mQED, which offers a phenomenological Green’s function quantization scheme for electromagnetic fields in the presence of spatially inhomogeneous, lossy and dispersive media gruner1996mqed; dung1998mqed; scheel1998qed; dung2000mqed; raabe2007unified, later generalized for systems with gain matloob1997electromagnetic; raabe2008qed. However, in order to facilitate a rigorous application of the theory, a brief discussion concerning some important prerequisites is warranted. The complex permittivity $\varepsilon(\bm{r},\omega)=\varepsilon_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\varepsilon_{\mathrm{i}}(\bm{r},\omega)$ and inverse permeability $\mu^{-1}(\bm{r},\omega)=\kappa(\bm{r},\omega)=\kappa_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\kappa_{\mathrm{i}}(\bm{r},\omega)$ are required to satisfy the Kramers-Kronig relations (i.e., they are analytic in the closed, upper-half, complex $\omega$ plane and $\zeta(\bm{r},\omega)\rightarrow 1$ for $|\omega|\rightarrow\infty$ ($\zeta\in\{\varepsilon,\kappa\}$). This is not really a requirement of the quantum theory, but arises from the physical necessity of causality. Furthermore, they should be reciprocal, $\zeta(\bm{r},\omega)=\zeta^{\mathrm{T}}(\bm{r},\omega)$, and satisfy the reflection symmetry condition $\zeta_{ij}(\bm{r},-\omega^{*})=\zeta^{*}_{ij}(\bm{r},\omega)$, which guarantees their reality in the time domain. These properties ensure that the fundamental QED commutation relations between the electric and magnetic field operators are maintained. If the material parameters $\zeta(\bm{r},\omega)$ of the system are reciprocal, it readily follows from Eq. (10) that the PML-transformed parameters $\tilde{\zeta}(\bm{r},\omega)$ are as well. It should be noted that there have been generalizations of mQED for the cases of non-reciprocal materials and materials with magnetoelectric crossing effects buhmann2012macroscopic, which are, however, beyond the scope of this work.
@@
When incorporating dispersive or non-dispersive PMLs in the quantization, two problems arise. For dispersive PMLs, it is easy to show from Eq. (10) and the Jacobian in Eq. (9) that the PML-transformed parameters $\tilde{\zeta}(\bm{r},\omega)$ satisfy reflection symmetry, which preserves the important property that modes with $\Re(\tilde{\omega}_{n})\neq 0$ always appear in pairs $(\tilde{\bm{E}}_{\mu},\tilde{\omega}_{\mu})$, $(\tilde{\bm{E}}^{*}_{n},-\tilde{\omega}^{*}_{n})$ and that all the modes have a non-positive imaginary part $\Im(\tilde{\omega}_{n})\leq 0$. Due to the $\omega$-dependence of the coordinate transformation, the material parameters have a pole at $\omega=0$ and consequently do not satisfy the Kramers-Kronig relations (i.e., they do not correspond to a causal medium). To resolve this issue, we could construct causal PMLs kuzuoglu2002frequency by making the replacement $\omega\rightarrow\omega+\mathrm{i}\eta$ in Eq. (6) with a small $\eta>0$, thereby shifting the pole into the lower, complex $\omega$ half-plane. Note that these causal PMLs also work, in principle, in the static case $\omega=0$, due to the finite values of the transformed material parameters at this point kuzuoglu2002frequency, whereas the usual PMLs fail and cannot be used to reproduce static modes besbes2022role. However, we are only interested in the fields in the inner, unmapped domain $\Omega\setminus\Omega_{\mathrm{PML}}$, and by using the properties of the Green’s function we will show that the correct commutator is maintained within this region even if it is not in the PML region $\Omega_{\mathrm{PML}}$, so we expect that the use of PMLs as defined in Eq. (6) should suffice in most cases where static modes can be neglected. When using non-dispersive PMLs, the PML-transformed parameters no longer satisfy reflection symmetry, and since they do not afford the damping of waves with non-positive frequencies, the QNMs with $\Re(\tilde{\omega}_{n})\leq 0$ are not correctly recovered and PML modes with $\Im(\tilde{\omega}_{n})>0$ occur, which would grow exponentially in time sauvan2022norm. In practice, one is often interested in the spectral response of a system only within a finite frequency interval $\Delta\omega$, in which case non-dispersive PMLs could be used provided that the contributions of the modes with $\Re(\tilde{\omega}_{n})\leq 0$ are negligible in $\Delta\omega$. With the same argument, one can also justify the use of material parameters $\zeta(\bm{r},\omega)$ or PMLs that do not satisfy the Kramers-Kronig relations (e.g., non-dispersive and complex), provided that the true, physical parameters are sufficiently well approximated within $\Delta\omega$.
@@
$\nabla\times\tilde{\kappa}\nabla\times G(\bm{r},\bm{r}^{\prime},\omega)-\frac{\omega^{2}}{c^{2}}\tilde{\varepsilon}G(\bm{r},\bm{r}^{\prime},\omega)=\bm{I}\delta(\bm{r}-\bm{r}^{\prime})\,.$ (18)
If the system exhibits loss only, then the Green’s function is analytic in the upper $\omega$ half-plane and has a set of discrete poles (corresponding to the QNMs), as well as possible branch cuts in the lower half-plane in the case of non-trivial background geometries. The introduction of gain components in the system shifts up the poles in the complex $\omega$ plane and a prerequisite to apply the theory for such active systems is that all the poles remain located in the lower half-plane. Again, this is not so much a fundamental limitation of the quantum theory but is, rather, a constraint on its range of validity for systems with linear amplification raabe2008qed. The Green’s function thus also satisfies the Kramers-Kronig relations, and besides that, reflection symmetry and the reciprocity $G(\bm{r},\bm{r}^{\prime},\omega)=G^{\mathrm{T}}(\bm{r}^{\prime},\bm{r},\omega)$.
@@
$\nabla\times\tilde{\kappa}\nabla\times\hat{\bm{E}}(\bm{r},\omega)-\frac{\omega^{2}}{c^{2}}\tilde{\varepsilon}\hat{\bm{E}}(\bm{r},\omega)=i\omega\mu_{0}\hat{\bm{j}}_{\mathrm{F}}(\bm{r},\omega)\,.$ (20)
@@
$+\nabla\times\bigg(\sqrt{-\frac{\hbar}{\pi\mu_{0}}\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega)}\hat{f}_{\mathrm{m},j}(\bm{r},\omega)\bm{e}_{j}\bigg)\,$,
in which we have used the Einstein summation convention and $\hat{f}_{\lambda,j}(\bm{r},\omega)$ ($\hat{f}^{\dagger}_{\lambda,j}(\bm{r},\omega)$) are bosonic annihilation (creation) operators of the joint (polaritonic) excitations of the field-dressed medium. The latter satisfy the commutation relationdung1998mqed
@@
The noise current defined in Eq. (21) stands in accordance with the fluctuation-dissipation theorem and ensures preservation of the fundamental commutation relations of the electric and magnetic fields. These formulae can also be derived from the canonical quantization of a microscopic model in which the electric and magnetic material responses are described using separate oscillator fields philbin2010canonical. From this perspective, it is clear that the Fock space corresponding to these bosonic variables describes not only the electromagnetic degrees of freedom, but also those of the material excitations.
@@
$+\nabla\times\bigg(\sqrt{\frac{\hbar}{\pi\mu_{0}}|\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega)|}\bm{e}_{j}\bigg[\Theta(-\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega))\hat{f}_{\mathrm{m},j}(\bm{r},\omega)$
@@
$\cdot\bigg($ $\mathrm{sgn}(\tilde{\varepsilon}_{\mathrm{i},j}(\bm{r},\omega))\hat{f}_{\mathrm{e},j}^{\dagger}(\bm{r},\omega)\hat{f}_{\mathrm{e},j}(\bm{r},\omega)$
@@
Note that excitations in the gain regions have a negative-energy contribution to the Hamiltonian. These expressions can again be derived from a microscopic model using inverted oscillator fields in the gain regions amooghorban2011casimir. However, it is worth underlining that the expressions in Eqs. (21) and (III) cannot be derived from a generalized phenomenological quantization scheme that allows for general non-local material responses described by a conductivity tensor, whenever the permittivity and permeability are spatially inhomogeneous raabe2008qed; raabe2007unified. Since these expressions can be motivated from a microscopic model and provide analytic results for arbitrary magnetodielectric media, we will make use of them in this paper.
@@
$=\frac{\hbar\mu_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\int_{\Omega}\mathrm{d}^{3}s$
@@
$-(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg]$
@@
where the latter equation shows that the commutator is not affected by the PMLs for $\bm{r},\bm{r}^{\prime}\in\Omega\setminus\Omega_{\mathrm{PML}}$. Note that the identity in Eq. (III) can only be used if the spatial integration is performed over the entire space $\Omega$ including the PMLs, so that the presence of the PMLs must necessarily be taken into account in the quantization.
By invoking the first equality and the Green’s function expansion in Eq. (17), we can expand the commutator in terms of the modes:
@@
$P_{nn^{\prime}}=\frac{\hbar\mu_{0}}{\pi}$ $\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\int_{\Omega}\mathrm{d}^{3}s\,A_{n}(\omega)A_{n^{\prime}}^{*}(\omega)$ (29)
$\cdot\bigg[$ $\frac{\omega^{2}}{c^{2}}\tilde{\bm{E}}_{n}(\bm{s})\tilde{\varepsilon_{\mathrm{i}}}(\bm{s},\omega)\tilde{\bm{E}}^{*}_{n^{\prime}}(\bm{s})$
$-(\nabla\times\tilde{\bm{E}}_{n}(\bm{s}))\tilde{\kappa}_{\mathrm{i}}(\bm{s},\omega)(\nabla\times\tilde{\bm{E}}^{*}_{n^{\prime}}(\bm{s}))\bigg]\,.$
@@
in which the modal operators can be expressed through the continuous bosonic operators:
$\hat{\xi}^{(\dagger)}_{n,\mathrm{e},\mathrm{L/G}}=\mathrm{i}\mu_{0}\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}A_{n}(\omega)\int_{\Omega}\mathrm{d}^{3}r\,\theta(\pm\tilde{\varepsilon}_{\mathrm{i},j}(\bm{r},\omega))$
@@
$\hat{\xi}^{(\dagger)}_{n,\mathrm{m},\mathrm{L/G}}=\mathrm{i}\mu_{0}\int_{0}^{\infty}\mathrm{d}\omega\,\omega A_{n}(\omega)\int_{\Omega}\mathrm{d}^{3}r\,\theta(\mp\tilde{\kappa}_{\mathrm{i},j}(\bm{r},\omega))$
@@
Note that the hermitian conjugates of the gain operators $\hat{\xi}_{n,\mathrm{e/m},\mathrm{G}}$ have to be used in the expansion of the positive-frequency part of the electric field, in order to reproduce the correct commutation relation as given in Eq. (28). Although they are hermitian conjugates of modal operators, they still behave like superpositions of positive-frequency excitations ($\propto\mathrm{e}^{-\mathrm{i}\omega t}$), as can be seen from Eqs. (III) and (32), and the Heisenberg equations for the continuum operators $\hat{f}_{\mathrm{e/m},j}(\bm{r},\omega)$ in the gain regions with the Hamiltonian of Eq. (25). This is analogous to the approach in Ref. franke2022coupled. In our case, however, we have included the PMLs into the quantization, which always exhibit gain. In the presence of gain materials, the vacuum correlation function of the electric field would be given by
@@
$=\frac{\hbar\mu_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\int_{\Omega}\mathrm{d}^{3}s$
@@
$+(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}|\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)|(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\bigg]\ .$
Due to the moduli of the material parameters in this formula, the identity in Eq. (III) cannot be used to express the correlation function in terms of the imaginary part of the Green’s function, which must be taken into account when, for example, calculating QE decay rates franke2021fermi or Casimir(-Polder) forces sambale2010casimir. At this point, it is also important to be aware of the fact that PMLs in 3D always possess gain components (i.e., not all of the components of the effective material parameters $\tilde{\varepsilon},\tilde{\mu}$ have a positive imaginary part). In our approach, the presence of these gain components in the PMLs would impact the correlation function due to induced fluctuations and thus the quantum dynamics, as they would, for example, affect the spontaneous decay of a QE. To see this, we can write the vacuum correlation function in our PML-transformed system as $\langle\hat{E}_{i}(\bm{r})\hat{E}_{j}(\bm{r}^{\prime})\rangle=(\hbar\mu_{0}/\pi)\int\mathrm{d}\omega\,\omega^{2}\mathrm{Im}(G_{ij}(\bm{r},\bm{r}^{\prime},\omega))+C_{\mathrm{gain}}^{\mathrm{res}}(\bm{r},\bm{r}^{\prime})+C_{\mathrm{gain}}^{\mathrm{PML}}(\bm{r},\bm{r}^{\prime})$, with
$C_{\mathrm{gain}}^{\mathrm{res/PML}}(\bm{r},\bm{r}^{\prime})=\frac{2\hbar\mu_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}\omega\,\omega^{2}\int_{\Omega{\mathrm{res/PML}}}\mathrm{d}^{3}s\,$
$\theta(-\tilde{\varepsilon}_{\mathrm{i},m}(\bm{s},\omega))\frac{\omega^{2}}{c^{2}}G^{\dagger}_{im}(\bm{s},\bm{r},\omega)|\tilde{\varepsilon}_{\mathrm{i},m}(\bm{s},\omega)|G_{mj}(\bm{s},\bm{r}^{\prime},\omega)$
$+\theta(\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega))(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}|\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)|$
$\cdot(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}\ ,$ (34)
which are the additional fluctuations caused by the gain components of the resonator and the PMLs. The correlations based on the term $C_{\mathrm{gain}}^{\mathrm{PML}}(\bm{r},\bm{r}^{\prime})$ are not physical, and we must therefore formulate our modal theory in such a way that it reproduces the correct correlation function without this term.
To achieve this, we redefine the gain and loss operators in Eqs. (III) and (32) by including the gain components of the PML region $\Omega_{\mathrm{PML}}$ in the loss operators:
@@
$(\nabla\times\bm{E}_{n}(\bm{r}))_{j}(\nabla\times\bm{E}_{n^{\prime}}^{*}(\bm{r}))_{j}$
@@
$(\nabla\times\bm{E}_{n}(\bm{r}))_{j}(\nabla\times\bm{E}_{n^{\prime}}^{*}(\bm{r}))_{j}\bigg]$
@@
$(\nabla\times\bm{E}^{*}_{n}(\bm{r}))_{j}(\nabla\times\bm{E}_{n^{\prime}}(\bm{r}))_{j}$ (42)
All the other commutators vanish due to the bosonic commutation relations of the operators $\hat{\bm{f}}_{\lambda}(\bm{r},\omega)$ and the distinct loss/gain regions. Since the commutation relations of these modal operators are non-bosonic, we have to apply symmetrization transformations separately to the electric/magnetic and loss/gain components in order to ascertain bosonic modal operators franke2019quantization; lowdin1950non:
@@
[…]
@@
$+(\nabla\times G(\bm{s},\bm{r},\omega))^{\dagger}_{im}|\tilde{\kappa}_{\mathrm{i},m}(\bm{s},\omega)|(\nabla\times G(\bm{s},\bm{r^{\prime}},\omega))_{mj}$